|x|+|x+5|=5 find x.

|x|+|x+5|=5  find x.

Grade:10

2 Answers

Sunil Kumar
10 years ago
for x>0 mod x = x mod(x+5) = x+5 for -5
Sunil Kumar
10 years ago
for x> 0 we get the equation 2x+5 = 5 so x = 0 for -5